OpenStudy (anonymous):

Use the table below to answer this question: x -1 3 5 y 7 3 1 Find the average rate of change for the given function from x = -1 to x = 5. Answers -6 -1 1 6

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Sorry i'm just not getting it,

OpenStudy (anonymous):

Ok basically the question is asking to find the average rate of change. This is confusing because of the way I listed the x and y values. Here is an image to help with that

OpenStudy (anonymous):

is it understandable? @dpasingh

OpenStudy (anonymous):

yes no maybe

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Average Rate of Change of a function, this is the change in the y-value divided by the change in the x-value for two distinct points on the graph.

OpenStudy (anonymous):

It would be \[\huge \frac{7-1}{5-(-1)} = \frac{6}{6} = 1\] Hence 3rd option is correct

OpenStudy (anonymous):

Whats your answer.?

OpenStudy (anonymous):

-6 -1 1 6